# Tensor tree learns hidden relational structures in data to construct generative models

### v1

@@ Discussion
We let the tensor tree represent a wave function in the Born machine. The Born machine framework has the following advantages. First, in principle, it may be implemented as a real quantum circuit, qualitatively improving upon any classical generative modeling. Though we have not explored this possibility in the present article, this is one of the attractive directions of future research. Second, it guarantees the positivity of the weights, which is the basis of the proposed method. However, we may consider other alternatives, such as the steepest descent by differentiating with respect to the logarithm of tensor elements rather than the tensor elements themselves. In addition, improvement in the optimization details is a subject of future work.
Methods
Tree tensor network representation for generative modeling
The proposed approach is based on optimizing a general tree structure for generative modeling. In the Born machine framework, the model distribution $p_{\theta}(\textbf{x})$ is defined by the Born rule in (1) as the squared amplitude of the wave function of a Born machine. Then, we introduce a tensor tree to represent the amplitude of the wave function [9]. The tensor train and balanced tree used in [2] and [3], respectively, are special cases of the general tree structure. Based on a tree structure, we can efficiently calculate various exact contractions and marginal distributions using recursive procedures.
We can define a canonical form for a tensor tree [15] to determine the optimal truncation from local calculations. We first choose the root edge to construct the canonical form of a given tensor tree. Then, all the tensors become isometries by recursively applying singular value decomposition (SVD) starting from the leaves. Finally, we obtain the matrix $\Lambda$ on the root edge. Using SVD, the root’s position can be moved to any neighboring edge. Our method chooses the edge to be updated as the root. After modifying the local connection and updating the tensors around the root edge, we move the root position to the neighboring edge that has not been updated for the longest time.
The Born machine is practically advantageous because there is no need to restrict the parameters to ensure the positivity of the distribution function. This is because the squared amplitude of the wave function represents it. Hence, we can use SVD to precisely decompose the whole distribution function into a tensor tree with the canonical form.
Statistical estimation of mutual information
$I(A,B)=\sum_{(a,b)}p(a,b)\ln\frac{p(a,b)}{p(a)p(b)},$ (3)
where $a$ and $b$ are configurations on sub-systems $A$ and $B$, respectively, $p(a,b)=p_{\theta}(\textbf{x}=(a,b))$ is a joint distribution, and $p(a)=\sum_{b}p(a,b)$ and $p(b)=\sum_{a}p(a,b)$ are marginal distributions. We can efficiently compute the values of $p(a)$ and $p(b)$ utilizing a tensor tree representation of a joint distribution in the canonical form. However, estimating the mutual information in a tensor tree state is challenging due to the need to calculate the summation for all configurations $(a,b)$. Moreover, the number of configurations grows exponentially with the system size.
where ${\cal M}^{\prime}$ is the ensemble of data samples. The estimator (4) may be negative if the empirical data distribution $\pi_{\textrm{emp}}(a,b)$ differs from the joint distribution $p(a,b)$.
An alternative method involves estimating the average of $\ln[p(a,b)/p(a)p(b)]$ using Monte Carlo sampling. If a Born machine has a tree structure, it can efficiently calculate the marginal distribution of a random variable given the other random variable states. The conditional distribution is then used to sample random variables perfectly. Repeating these steps can generate an independent sample $(a,b)$ with probability $p(a,b)$. Consequently, we can efficiently do Monte Carlo estimation of (3) as (4) with ${\cal M}^{\prime}$, the ensemble of samples generated from $p(a,b)$ directly. However, during learning, we only used the first method due to the low quality of the Born machine at the early stage.
Branch-reconnection algorithm
The detail of our branch-reconnection in Fig.1(b) is as follows:
  1. (i)
    Initialize the network and all tensor elements.
  2. (ii)
  3. (iii)
  4. (iv)
  5. (v)
    Replace the 4-leg tensor with the chosen decomposition.
  6. (vi)
    If the termination condition is satisfied, terminate.
  7. (vii)
    Move to a bond connected to one of the two new tensors and return to (iii).
We utilized a reconnection-based decomposition of a composite tensor to calculate the ground state of a quantum model with a tensor tree[16]. We improve the tensor in (iii) and (iv) by employing gradient-descent updates to minimize NLL.
Training details
In steps (iii) and (iv) in the branch-reconnection algorithm, we improve the combined tensor and new tensors in the decomposition by employing gradient-descent updates to minimize NLL [2, 3]. Assuming that the target distribution of generative modeling is an empirical data distribution, $\pi_{\textrm{emp}}$, NLL is defined as
where ${\cal M}$ is the ensemble of data samples and $D_{\textrm{KL}}$ is a Kullback-Leibler divergence. In this study, ${\cal M}$ is the same as ${\cal M}^{\prime}$ in (4). We perform a single gradient-descent update in the procedure (iii) and ten ones in (iv), with learning rates of 0.05 for random patterns and 0.001 for others. For data samples ${\cal M}$, we use the batch of training data samples for random patterns and S&P 500 stock price fluctuation patterns and a mini-batch of size 1000 for others. If the mini-batch size is small, instability occurs in estimating mutual information. We set the interval for changing the mini-batch to 1000 gradient-descent updates for the handwritten digits and updated the new mini-batch after all edges for the Bayesian networks. The NLL for test data is useful for checking the overfitting of learning data. For the Bayesian network data, we showed the optimized network with lower NLL for test data in Fig.3 because it fluctuates in the optimization process when changing a mini-batch.
We use the QMNIST dataset[10] for the images of handwritten digits. It is an improved version of the Mixed National Institute of Standards and Technology database (MNIST) dataset[17], a collection of handwriting digits from 0 to 9 on $28\times 28$ pixels. In QMNIST, the number of test images increases to the number of learning images, 60000. The intensity of a pixel ranges from 0 to 255. As in [3], by padding zero intensity pixels around an original image, we embed an original image into $32\times 32$ pixels. We make a binary bit from a pixel’s intensity. If the intensity is larger than 127, the bit is set to one; otherwise, it is set to zero. The final image consists of $32^{2}$ bits.
In the Bayesian networks, each random variable is binary. We generate the target data according to the causal relation illustrated in Fig.3 (a-c). When there is a single cause, say $A$, we generate the result bit $B$ with the conditional probability $P(B=A|A)=r$. In contrast, when there are multiple causes, we first calculate the intermediate bit using the exclusive-or of all the cause bits and generate the result bit with the same conditional probability. The present article’s correlation rate $r$ is fixed to be 0.8. In all cases, we use a bond dimension of 4 for a tensor tree.
The pseudo-program is terminated at (vi) when the number of iterations reach a prefixed number, $N$. In the random patterns, handwritten digits, Bayesian networks, and S&P 500 examples, we choose $N=2000,10^{6},3000$ and $2\times 10^{6}$, respectively.
The initial random tree structure influences the results. Therefore, we perform five optimizations to assess the variance of NLLs in Fig.2(b) and (e), and Fig.3(d). We generated an initial random tree as follows:
  1. (i)
  2. (ii)
  3. (iii)
    Until the list S contains only one element, repeat step (ii).
The final element in S is a tree with the random variables associated with its leaves.
@@ Acknowledgments
This work was partially supported by the joint project of Kyoto University and Toyota Motor Corporation, titled “Advanced Mathematical Science for Mobility Society.” K. H. appreciates fruitful comments from T. Hikihara, H. Ueda, K. Okunishi, T. Nishino, and L. Wang and acknowledges the support from JSPS KAKENHI (Grant No. 20K03766 and 24K06886) and a Grant-in-Aidfor Transformative Research Areas “The Natural Laws of Extreme Universe—A New Paradigm for Spacetime and Matter from Quantum Information” (KAKENHI Grants No. 21H05182 and No. 21H05191) from JSPS of Japan. T. O. acknowledges the support from JSPS KAKENHI (Grant No. 23H03818 and 22K18682), the Endowed Project for Quantum Software Research and Education, The University of Tokyo (https://qsw.phys.s.u-tokyo.ac.jp/), and The Center of Innovations for Sustainable Quantum AI (JST Grant No. JPMJPF2221). N. K. acknowledges the support from JSPS KAKENHI (Grant No. 23H01092). The computation in this work has been done using the facilities of the Supercomputer Center, the Institute for Solid State Physics, the University of Tokyo.

### v2

@@ 4 Discussion and Conclusion
We let the tensor tree represent a wave function in the Born machine. Theoretically, a tensor tree state can be implemented as a quantum circuit. An optimized tensor network was utilized for preparing the parametrized quantum circuits of the Born machine to resolve serious optimization issues of the quantum circuits, the prevalence of barren plateaus in optimization landscapes[14]. Though we have not explored this possibility in the present article, this is one of the attractive directions of future research.
In summary, the proposed method optimizes a tree structure to minimize BMI, enhancing performance and revealing hidden relational structures in the target data.
Data availability statement
The data that support the findings of this study are openly available at the following URL/DOI: https://github.com/KenjiHarada/adaptive-tensor-tree-generative-modeling[30].
@@ Acknowledgments
This work was partially supported by the joint project of Kyoto University and Toyota Motor Corporation, titled “Advanced Mathematical Science for Mobility Society.”[31] K. H. appreciates fruitful comments from T. Hikihara, H. Ueda, K. Okunishi, T. Nishino, and L. Wang and acknowledges the support from JSPS KAKENHI (Grant No. 20K03766 and 24K06886) and a Grant-in-Aidfor Transformative Research Areas “The Natural Laws of Extreme Universe—A New Paradigm for Spacetime and Matter from Quantum Information” (KAKENHI Grants No. 21H05182 and No. 21H05191) from JSPS of Japan. T. O. acknowledges the support from JSPS KAKENHI (Grant No. 23H03818 and 22K18682), the Endowed Project for Quantum Software Research and Education, The University of Tokyo (https://qsw.phys.s.u-tokyo.ac.jp/), and The Center of Innovations for Sustainable Quantum AI (JST Grant No. JPMJPF2221). N. K. acknowledges the support from JSPS KAKENHI (Grant No. 23H01092). The computation in this work has been done using the facilities of the Supercomputer Center, the Institute for Solid State Physics, the University of Tokyo.